OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4.

OpenStudy (jackellyn):

Solve for y -4y=4-x y=(4-x)/-4 Simplify y=-1 +x/4 so Y intercept is -1 and the slope is 1/4